OpenStudy (anonymous):

Can someone help? I cannot find in my book how to do this and and I can't find any further information online. Given triangle QRS is congruent to triangle TUV, QS=3v+2 and TV=7v-6, find the length of QS and TV

hero (hero):

\[\triangle{QRS} ≅ \triangle{TUV} \\\text{Then} \space QS = TV \\\text{Thus} \space 3v+2 = 7v - 6 \] Solve for v

OpenStudy (anonymous):

okay, what would be the first step? Would you first add the 2 and take away the 6?

OpenStudy (anonymous):

or would you first try to get v on it's own?

OpenStudy (anonymous):

Get your variable numbers on one side, and the normal numbers on the other

OpenStudy (anonymous):

3v + 2 - 2 = 7v - 6 + 6

OpenStudy (anonymous):

3v + 7v = 7v - 7v + 6

OpenStudy (anonymous):

wait, messed that up

OpenStudy (anonymous):

Then you would have 3v=7v if you cancel out each side right? But how do you go from there?

hero (hero):

@seantay, don't confuse the poor guy

OpenStudy (anonymous):

@Hero I was taught that you only cancel out/ add or take away with one of the numbers not both? Is it supposed to be ending up with 3v=7v?

hero (hero):

3v + 2 = 7v - 6 2 + 6 = 7v - 3v 8 =4v 8/4 = v 2 = v Length of QS = 3(2) + 2 = 8 Length of TV = 7(2) - 6 = 14 - 6 = 8

OpenStudy (anonymous):

@Hero Thank you for your explanation, that matches what I have done in the past but I sometimes doubt if I am using the right way of solving a problem, I didn't know if there was a new way of solving this type of problem since it relates to congruence and angles.